Please help!!!
Find the difference quotient f(x+h)-f(x)over h for 2-x^2
THANKS!!!

Answer by stanbon(75887)

Find the difference quotient f(x+h)-f(x)over h for 2-x^2
--------------------
f(x+h) = 2 - (x+h)^2
f(x+h) = 2 -x^2-2hx-h^2
f(x) = 2 - x^2
--------------
f(x+h)-f(x) = -2hx-h^2
--------------
[f(x+h)-f(x)]/h = -2x-h
===============================
Cheers,
Stan H.